OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a.  4x2 - 32x + 16 = 0

myininaya (myininaya):

First of all doesn't 4 divide 4,32,and 16 So we can factor out 4 and divide it on both sides since 4 doesn't =0 \[4(x^2-8x+4 )=0\] \[\frac{4}{4} (x^2-8x+4)=\frac{0}{4}\] \[x^2-8x+4=0\] Now: Lets talk about completing the square.... \[x^2+bx+(\frac{b}{2})^2=(x+\frac{b}{2})^2\] So if b=-8, then what is \[(\frac{b}{2})^2 ?\]

OpenStudy (anonymous):

this is algebra your solving to find what c=

OpenStudy (anonymous):

4x^2-32x+16=0 taking 4 common 4(x^2-8x+4)=0 x^2-8x+4=0 x^2-8x+4^2=-4+4^2 (x-4)^2=12 x-4=squareroot12 x-4=+-2squareroot3 x-4=2squareroot3 x=2squareroot3+4 and x=-2squareroot3-4 are the answer